OpenStudy (anonymous):

find 1+2+3+4....+80 the sum of the 80 natural numbers the sum is ?

OpenStudy (campbell_st):

here is an old method reverse the order and sum the pairs 1 2 3 4 5 6 .... 79 80 + 80 79 78 77 76 75... 2 1 ------------------------------ what does each pair add 2

OpenStudy (anonymous):

i dont get it

OpenStudy (campbell_st):

ok... so what does each pair of numbers add to... of you add down each column

OpenStudy (anonymous):

ok thanks

OpenStudy (campbell_st):

lol... why worry about learning.... when people just give answers... good luck

OpenStudy (anonymous):

can u elavorate? i would like to learn too!

OpenStudy (campbell_st):

well each vertical pair adds to 81 does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then?

OpenStudy (campbell_st):

well there are 80 pairs of numbers and you have added the same number's twice... so the sum is Sum = 80 x 81/2

OpenStudy (campbell_st):

so in general, if you know the 1st and last number in an arithmetic sequence the sum is \[S_{n} = \frac{n \times(a + l)}{2}\] n = number of terms, a = 1st term and l = last term. if you don't know the last term you need to use \[S_{n} = \frac{n}{2} [ 2a + (n -1)\times d]\] d is the common difference in the sequence... or what it is always changing by. hope this make sense

OpenStudy (anonymous):

yes thanks as well